5+(-19)+(+2)-(-5)
1
Expert's answer
2012-05-08T07:31:57-0400
5+(-19)+(+2)-(-5)=5-19+2+5=-7
If there is "+" before braces, numbers remain the same after opening. if there is "-" before braces when opening we change the sign of every number